A certain type of storage battery lasts on average, 3.0 years with a standard deviation of 0.5 year. Assuming the battery life is normally distributed, find the probability that a given battery will last more than 2 years.


1
Expert's answer
2021-11-12T13:46:28-0500

"\mu=3.0 \\
\sigma= 0.5 \\
P(X>2) = 1 -P(X<2) \\
= 1 -P(Z< \frac{2-3}{0.5}) \\
= 1 -P(Z< -2) \\
= 1 -0.0227 \\
= 0.9773"